three years ago,Rica was thrice as old as Dandy. three years hence, she will be twice as old. find their present age.

Answer by swincher4391(1107)

3*(d-3) = (r-3)
2*(d+3) = (r+3)
3d - 9 = r-3
2d + 6 = r+3
------------
d - 15 = -6
d = 9
3(9) - 9 = r -3
18 = r -3
r = 21